# Can Helicity Modulus Be Defined For Boundary Conditions With Finite Twist?

### v1

@@ 4 Results
Figure 4: Excess entropy $\delta S$ due to the mixing of chiralities vs. temperature $T$ for different system sizes. The solid horizontal line corresponds to $\delta S=\ln(2)$. The inset shows the extrapolation of the maximum value of each $\delta S$ vs. $T$ curve to infinite system size. The extrapolation gives $\delta S_{max}(L=\infty)=0.69(1)$ and agrees with $\ln(2)$.
The figure shows that $\delta S$ approaches zero with increasing system size for high temperatures while it appears to approach a constant for low temperatures. In fact, the figure suggests that $\delta S$ approaches a step function in the thermodynamic limit. To determine the step height and position, we extrapolate the peak value $\delta S_{max}$ of the $\delta S$ vs. $T$ curves using,
$\delta S_{max}(L)=\delta S_{max}(\infty)+aL^{-\lambda}$ (9)
where $a,\lambda$ are fitting parameters. (We determine $\delta S_{max}$ as an extremum of a quadratic curve fitted in the vicinity of the peak.) The extrapolation is shown in the inset of Fig. 4 and yields $\delta S_{\infty}=0.69(1)\approx\ln(2)$, in agreement with the expectation of contributions from two degenerate states. Additionally, an extrapolation of the temperature at which the $\delta S$ vs. $T$ curves cross the $\delta S=\ln(2)$ line matches with $T_{c}$ within the error bars. Thus, our numerical data extrapolate to $\delta S=\ln(2)\Theta(T_{c}-T)$ where $\Theta$ is the Heavyside step function. We note that the arguments predicting the excess $\ln(2)$ entropy due to the chirality mixing do not include states with higher winding numbers. These states are known to renormalize the helicity modulus in two dimensions [20, 19], but the effect is tiny and only visible in high-accuracy simulations beyond our numerical precision.
The excess entropy due to the chirality mixing reduces the free energy cost of a $\pi$ twist by $T\ln(2)$. Consequently, the helicity modulus $\rho_{s}(\pi)$ computed directly from eq. (3) is reduced by $2\ln(2)T/\pi^{2}$. This explains the observation in Fig. 1 of a lower helicity modulus for the $\pi$ twist. Consequently, if one wishes find the critical temperature from simulations employing a $\pi$ twist (i.e., anti-periodic boundary conditions), the reduction of $\rho_{s}$ has to be accounted for. This can be achieved by modifying the Kosterlitz-Nelson relation by changing the slope of the line crossing the $\rho_{s}$ curves from $2/\pi$ to $2/\pi-2~{}\ln(2)/\pi^{2}$. The resulting analysis is presented in Fig. 5 which shows the helicity modulus for a $\pi$ twist as a function of temperature for different system sizes.
Figure 5: Helicity modulus $\rho_{s}$ for a twist of $\Theta=\pi$ (anti-periodic boundary conditions). The inset shows the intersection temperatures $T^{*}$ of the helicity modulus curves with the modified Kosterlitz-Nelson line of slope $2/\pi-2~{}\ln(2)/\pi^{2}$. An extrapolation according to eq. (7) yields $T_{c}=0.897(2)$.
@@
5 Conclusion
@@
To summarize, in this paper we have studied the effects of finite (non-infinitesimal) twists in the boundary conditions on a two-dimensional classical ferromagnetic XY model. The system’s response to the twist has been studied by direct evaluation of the free energy by means of Monte Carlo simulations. We have found that in the quasi long-range ordered phase, the free energy cost of a non-infinitesimal twist deviates from the expected quadratic dependence on the twist angle. In the case of a $\pi$ twist (anti-periodic boundary conditions), the mixing of states of opposite chiralities causes an excess entropy contribution of $\ln(2)$ which reduces the observed helicity modulus compared to the value for an infinitesimal twist. This analysis has allowed us to accurately determine the helicity modulus from simulations with anti-periodic boundary conditions. A modified Kosterlitz-Nelson relation can be used to determine the critical temperature.
@@
It is also interesting to ask what implications our findings have in the context of the quantum-to-classical mapping. The partition function of a quantum system in $d$ dimensions can often be mapped onto that of a classical system in $d+1$ dimensions. How does the extra entropy contribution in the classical Hamiltonian with $\pi$-twisted boundary conditions manifest itself in the ground state of the corresponding quantum system?
6 Acknowledgements
The authors acknowledge support from the National Science Foundation under grant nos. DMR-1506152, DMR-1828489, OAC-1919789. Simulations were performed on the Foundry and Pegasus Clusters at Missouri University of Science and Technology.

### v3

@@ IV.1 Free energy response to finite twist
Figure 4: Excess entropy $\delta S$ due to the mixing of chiralities vs. temperature $T$ for different system sizes. The solid horizontal line corresponds to $\delta S=\ln(2)$. The inset shows the extrapolation of the maximum value of each $\delta S$ vs. $T$ curve to infinite system size. The extrapolation gives $\delta S_{\textrm{max}}(L=\infty)=0.69(1)$ and agrees with $\ln(2)$.
The figure shows that $\delta S$ approaches zero with increasing system size for high temperatures while it appears to approach a constant for low temperatures. In fact, the figure suggests that $\delta S$ approaches a step function in the thermodynamic limit. To determine the step height and position, we extrapolate the peak value $\delta S_{\textrm{max}}$ of the $\delta S$ vs. $T$ curves using,
$\delta S_{\textrm{max}}(L)=\delta S_{\textrm{max}}(\infty)+aL^{-\lambda}$ (10)
where $a,\lambda$ are fitting parameters. (We determine $\delta S_{\textrm{max}}$ as an extremum of a quadratic curve fitted in the vicinity of the peak.) The extrapolation is shown in the inset of Fig. 4 and yields $\delta S_{\infty}=0.69(1)\approx\ln(2)$, in agreement with the expectation of contributions from two degenerate states. Additionally, an extrapolation of the temperature at which the $\delta S$ vs. $T$ curves cross the $\delta S=\ln(2)$ line matches with $T_{c}$ within the error bars. Thus, our numerical data extrapolate to $\delta S=\ln(2)\Theta(T_{c}-T)$ where $\Theta$ is the Heavyside step function. We note that the arguments predicting the excess $\ln(2)$ entropy due to the chirality mixing do not include states with higher winding numbers. These states are known to renormalize the helicity modulus in two dimensions [37, 32], but the effect is tiny and only visible in high-accuracy simulations beyond our numerical precision.
The excess entropy due to the chirality mixing reduces the free energy cost of a $\pi$ twist by $T\ln(2)$. Consequently, the helicity modulus $\rho_{s}(\pi)$ computed from eq. (3) is reduced by $2\ln(2)T/\pi^{2}$ compared to the infinitesimal twist value arising from eq. (4). This explains the observation in Fig. 1 of a lower helicity modulus for the $\pi$ twist. Consequently, if one wishes find the critical temperature from simulations employing a $\pi$ twist (i.e., anti-periodic boundary conditions), the reduction of $\rho_{s}$ has to be accounted for. This can be achieved by adding the correction to the Kosterlitz-Nelson relation by changing the slope of the line crossing the $\rho_{s}$ curves from $2/\pi$ to $2/\pi-2~{}\ln(2)/\pi^{2}$. The resulting analysis is presented in Fig. 5 which shows the helicity modulus for a $\pi$ twist as a function of temperature for different system sizes.
Figure 5: Helicity modulus $\rho_{s}(\pi)$ for a twist of $\Theta=\pi$ (anti-periodic boundary conditions). The inset shows the intersection temperatures $T^{*}$ of the helicity modulus curves with the modified Kosterlitz-Nelson line of slope $2/\pi-2~{}\ln(2)/\pi^{2}$. An extrapolation according to eq. (7) yields $T_{c}=0.897(2)$.
@@
IV.2 Breaking of the chiral symmetry
In Sec. IV.1, we have established that, for an imposed twist $\Theta=\pi$, the thermodynamic ensemble contains an equal mixture of states with clockwise and counter-clockwise local twists. Now we explore the question whether a single given macroscopic system fluctuates between clockwise and counter-clockwise orientations or whether it spontaneously breaks the chiral symmetry below $T_{c}$ by freezing into one orientation. As the classical XY Hamiltonian does not contain any dynamic terms, the answer to this question will depend on the assumed dynamics of the system. Here, we focus on dissipative local model-A dynamics according to the Hohenberg-Halperin classification [38], but the results should hold qualitatively for other local dynamics as well.
To analyze chiral fluctuations, we define the bulk chirality $h$ via the vector product between nearest neighbor spins along the direction of the twist. Assuming a twist in the $x$-direction, $h$ is given by
$h=\frac{1}{L}\sum_{\langle ij\rangle_{x}}\mathbf{\hat{k}}\cdot\left({\mathbf{S}}_{i}\times{\mathbf{S}}_{j}\right)$ (11)
where site $j$ is the nearest neighbor of site $i$ in the positive $x$ direction, and $\mathbf{\hat{k}}$ is the unit vector in the $z$-direction. This means positive $h$ correspond to a counter-clockwise twist and negative $h$ correspond to a clockwise twist.
We monitor the time evolution of the bulk chirality $h$ in long Monte Carlo runs that perform only local Metropolis updates (implementing model-A dynamics). Specifically, we measure the rate $\Lambda$ at which the chirality changes sign during such a simulation. Figure 6 shows the flip rate $\Lambda$ as a function of temperature for different system sizes observed for runs of $1.2\times 10^{6}$ Metropolis sweeps averaged over $400$ samples.
Figure 6: Flip rate $\Lambda$ of the bulk chirality $h$ as a function of the inverse temperature $1/T$ for different system sizes $L$. The solid lines are fits to the ansatz (12). The resulting system size dependence of the activation energy $\Delta E^{\mathrm{flip}}$ is shown in the inset in a semi-logarithmic plot. The dotted line is a fit to the function, $\Delta E^{\mathrm{flip}}=E_{0}\ln(k_{0}L)$, giving fit parameters $E_{0}=7.91(6)$ and $k_{0}=0.39(1)$.
The data clearly suggest an exponential dependence of the flip rate on the inverse temperature. Indeed, if chirality flips are governed by activation over an energy barrier $\Delta E^{\mathrm{flip}}$, the flip rate is expected to follow the ansatz
$\Lambda=cT^{b}\exp(-\Delta E^{\mathrm{flip}}/T)~{}.$ (12)
where $b$ and $c$ are fit parameters. Figure 6 demonstrates that the observed flip rates indeed follow this ansatz within their statistical errors. The activation energies $\Delta E^{\mathrm{flip}}$ obtained from the fits in Fig. 6 are presented in the inset as a function of the system size in a semi-logarithmic plot. The data are well described by the logarithmic function
$\Delta E^{\mathrm{flip}}=E_{0}\ln(k_{0}L)$ (13)
with $E_{0}=7.91(6)$ and $k_{0}=0.39(1)$. This logarithmic dependence suggests that flips of the bulk chirality are facilitated by a vortex mechanism.
Figure 7 illustrates how the creation and annihilation of a vortex-antivortex pair can reverse the sign of the bulk chirality.
Figure 7: Schematic illustrating how a vortex-antivortex pair can flip the bulk chirality. For details see text.
Imagine a system with fixed boundaries in the $x$-direction that impose a $\pi$ twist and periodic boundary conditions in the $y$-direction. Initially, the system features a uniform clockwise twist to satisfy the boundary conditions. When a vortex-antivortex pair is introduced as shown in the figure, the spins between the vortices now rotate in the opposite direction. Now imagine that the + vortex travels upwards across the (periodic) boundary, approaches the -- vortex from below, and finally annihilates. The end result is a state with a uniform counter-clockwise twist, i.e., the chirality $h$ has changed sign. The activation energy of this process is given by the energy of a vortex-antivortex pair at the largest distance $L/2$. In the continuum limit, the energy of a vortex-antivortex pair at distance $D$ is given by $E_{\mathrm{pair}}=E_{\mathrm{core}}+2\pi J\ln\left(D/a)\right)$ where $E_{\mathrm{core}}$ is the vortex core energy and $a$ the core size. Neglecting the core energy (which can be formally absorbed by a shift of $a$), a comparison with eq. (13) gives a reasonable agreement. This indicates that the leading contribution to $\Delta E^{\mathrm{flip}}$ comes from the formation of a vortex-antivortex pair. Deviations can be attributed to finite-size effects, contributions from other flip processes and uncertainties in the sequence of fits leading to eq. (13)  [^2] If one considers open rather than periodic boundary conditions in the direction perpendicular to the twist, a single vortex moving across the system is sufficient to flip the bulk chirality. Test simulations with open boundary conditions lead to the same logarithmic dependence (13) of the activation energy $\Delta E^{\mathrm{flip}}$ but with prefactor $E_{0}=3.7(1)$, indicating the formation of a single vortex..
The results of this section demonstrate that the energy required to flip the bulk chirality diverges with system size in the thermodynamic limit. The divergence is only logarithmic in $L$ in contrast to the case of a domain wall whose energy would diverge linearly in $L$ in two dimensions. Nonetheless, it implies that a macroscopic system will not fluctuate between states of opposite bulk chirality during its time evolution (for local dynamics). In other words, the chiral symmetry is the spontaneously broken below the BKT transition.
V Conclusion
@@
To summarize, in this paper we have studied the effects of finite (non-infinitesimal) twists in the boundary conditions on a two-dimensional classical ferromagnetic XY model. The system’s response to the twist has been studied by a direct evaluation of the free energy by means of large-scale Monte Carlo simulations. We have found that in the quasi long-range ordered phase below the BKT transition, the free energy cost of a non-infinitesimal twist deviates from the expected quadratic dependence on the twist angle. In the case of a $\pi$ twist (anti-periodic boundary conditions), the mixing of states of opposite chiralities causes an excess entropy contribution of $\ln(2)$ that lowers the free energy cost of the twist. Thus, if the helicity modulus is calculated from the free energy of a $\pi$-twisted system (using eq. (4)), its value is reduced compared to the usual helicity modulus obtained from an infinitesimal twist. We note in passing that our discussion has been based on the phase representation of the XY model, but the same results are expected in other representations [40, 41, 42, 43] provided they correctly reflect the chiral symmetry of the boundary conditions.
A macroscopic $\pi$-twisted system in the quasi long-range ordered phase spontaneously breaks the chiral symmetry between states with clockwise and counter-clockwise local twists (at least in the case of local dynamics). This implies that experiments on a single macroscopic system cannot observe the $\ln(2)$ entropy contribution from chirality mixing and the corresponding reduction of the apparent helicity modulus value. The excess entropy is expected to be observable in mesoscopic systems that are small enough to fluctuate between states with opposite chirality and in situations involving not just a single system but an entire (thermodynamic) ensemble.
Chirality mixing is important for computer simulations for at least two reasons. First, efficient simulations often involve nonlocal algorithms (such as the Wolff cluster algorithm) that can freely flip the bulk chirality of the system. Moreover, system sizes in computer simulations are sometimes not very large so that even algorithms with local updates (such as the Metropolis algorithm) may allow the system to fluctuate between different chiralities. Free energies of $\pi$-twisted systems computed in such simulations contain the $-T\ln(2)$ entropic free energy contribution due to chirality mixing which needs to corrected to compute the usual (infinitesimal-twist) helicity modulus $\rho_{s0}$ from simulations with a $\pi$ twist. When determining the BKT transition temperature from the helicity modulus data, one can, alternatively, modify the Kosterlitz-Nelson relation as discussed in Sec. IV.1.
The partition function of a quantum system in $d$ dimensions can often be mapped onto that of a classical system in $d+1$ dimensions. Our findings imply that a macroscopic one-dimensional quantum rotor model with antiperiodic boundary conditions has a doubly degenerate ground state because the tunneling time between states of opposite chirality diverges with system size.
@@
Acknowledgements.
We acknowledge support from the National Science Foundation under grant nos. DMR-1506152, DMR-1828489, OAC-1919789. Simulations were performed on the Foundry and Pegasus Clusters at Missouri University of Science and Technology. We would also like to thank John Chalker, Eduardo Fradkin, and Martin Hasenbusch for helpful correspondence and W. Joe Meese for valuable discussions.